# Toward building next-generation Geocoding systems: a systematic review

### v1

@@ Abstract
Geocoding systems are widely used in both scientific research for spatial analysis and everyday life through location-based services. The quality of geocoded data significantly impacts subsequent processes and applications, underscoring the need for next-generation systems. In response to this demand, this review first examines the evolving requirements for geocoding inputs and outputs across various scenarios these systems must address. It then provides a detailed analysis of how to construct such systems by breaking them down into key functional components and reviewing a broad spectrum of existing approaches, from traditional rule-based methods to advanced techniques in information retrieval, natural language processing, and large language models. Finally, we identify opportunities to improve next-generation geocoding systems in light of recent technological advances.
@@ 1 Introduction
The geocoding process converts human-readable location descriptions into machine-readable geographic coordinates [34, 4]. The earliest known example dates back over two centuries, when a British doctor employed a dot map to visualize cholera cases [71]. In recent decades, geocoding has expanded beyond academic spatial analysis into widespread everyday applications [109, 75, 85, 86, 61, 110, 11]. This expansion has been driven by improved match rates and spatial accuracy resulting from better reference datasets, enhanced address models, and diverse data sources [2, 99, 114], along with rapid technological advances in web services and GPS-enabled mobile devices. Consequently, geocoding services have become readily accessible for daily use, and spatial analyses increasingly depend upon accurate geocoded data [86, 75].
Significant research efforts in academia and industry have targeted various aspects of the geocoding workflow, including address parsing [14, 74, 62], feature matching techniques [81, 65, 30, 76], interpolation methods [29, 7, 3], candidate ranking procedures [32, 54], and the establishment of comprehensive evaluation frameworks [36, 13]. Additional studies have addressed the impact of geocoded data quality on subsequent spatial analyses [116, 36]. Research on reverse geocoding has specifically emphasized reducing query latency [77, 68], integrating richer reference datasets [91], and improving the quality of output candidates [43, 106, 112, 9]. Previous review articles have also explored geocoding from various perspectives, including error sources and ongoing challenges [34], its implications for health-related research [86], and applications within Location-Based Services (LBS) alongside considerations for privacy protection [92].
@@
To inspire future advancements in geocoding techniques, this article begins by analyzing current demands for geocoding input and output. We then systematically decompose a geocoding system into key components based on their roles in the workflow, examining their underlying mechanisms and existing approaches. Finally, we identify research directions for the development of next-generation geocoding systems. This approach-oriented and component-based perspective sets our review apart from the prior literature. The remainder of this paper is organized as follows: Section 2 summarizes the variations in geocoding inputs, while Section 3 outlines the demands for geocoding outputs. Section 4 details the essential components for constructing geocoding systems. Section 5 identifies key opportunities for advancing geocoding techniques. Finally, Section 6 concludes with a summary of this survey.
2 Geocoding Input
Geocoding input reflects human understanding and description of locations. Table 1 summarizes several distinct location descriptions that have been used as geocoding input for various purposes and studies.
@@ 2 Geocoding Input
Table 1: Location descriptions for different application scenarios
@@
Traffic accidents [58] Distance/Topology related to intersections or landmarks 50 meters to the East of the intersection of Texas Ave. and Wellborn Rd.
Disease Outbreaks [123] Residential Postal Address 209 Augsburg Ct., College Station TX 77843
Crime Analysis [75]
Exposure Risk Assessment [86]
Location Based Services [95] Point of Interests (POI) Descriptions, Distance/Topology related to POIs Eiffel Tower, Downtown Paris
@@
Named Entity Recognition (NER) from Articles [83] Administration Regions (e.g., city, state) The Statue of Liberty is in New York Harbor in New York City, in the United States.
Indoor Navigation [24] Room Identifications + Building Identifications Room SW160, Building E
@@
These variations can be explained by three factors. First, geocoding input describes where events or incidents often occur. Case reporting systems typically regulate the format of location descriptions during a data collection process. For instance, patients are required to provide their residential postal addresses upon hospital admission. Similarly, in Texas, the car crash report instructions  [^1] https://ftp.dot.state.tx.us/pub/txdot-info/trf/crash_notifications/2023/cr100-v26.1.pdf mandate that each crash location be described in relation to the nearest intersections, referenced landmarks, or highway segments. Second, geocoding input reflects human cognition of locations [47]. For example, the region termed “downtown” does not physically exist on the earth but is defined by human activities, functions, and semantics [28]. Such semantic descriptions frequently appear in daily communication, often containing abbreviations and aliases for locations [31]. With the increasing prevalence of LBS such as ride-sharing and navigation in daily life [1, 95], geocoding inputs now range from formatted postal addresses to vernacular points of interest (POI) descriptions [72, 31]. Third, geocoding inputs stem from how humans artificially encode spatial areas. Different encoding schemes use combinations of alphabets and digits to represent spatial partitions, which ultimately serve as geocoding inputs. The most common encoding method is postal codes, assigned based on regional delivery needs [18, 50]. Similarly, room identification in buildings—encoding floor levels and segmentations based on arbitrary rules—functions as an input for indoor navigation scenarios.
@@
These variations in input format and content pose significant challenges for geocoding systems. Moreover, three key linguistic complexities further complicate the geocoding process. The first challenge is the semantics and ambiguity of words. Similar to semantic issues, ambiguity is a well-known problem in geocoding and Geographic Information Retrieval (GIR) domains [46, 31]. The ambiguityuity between street direction prefixes and street names frequently appears in geocoding practice. For example, ”N Main St.” could refer to a street named ”North Main St.” or ”Main St.” with a prefix of ”North.” The second challenge is that geocoding input is error-prone. A study has categorized input errors into two types: semantic errors and syntax errors [51]. Semantic errors are the result of the substitution of words with similar meanings. For example, a user might input ”Little River Rd” instead of ”Small Creek Rd”. Despite their similar meanings, geocoding systems struggle to distinguish them based solely on pronunciation or string similarities. Syntax errors often stem from habitual typing habits. For instance, the city name or the state name is often omitted in an address description [14]. This omission can affect the reference feature matching process, as the same street name can be found in many cities. Besides missing address elements, typographic errors frequently occur in geocoding input and are found to be sensitive to keyboard layout [56, 39]. Studies have identified the most frequent typographic error patterns in geocoding input as follows [13, 42, 81, 14, 21]: (1) replacing one character, (2) deleting one character, (3) inserting one character, and (4) swapping two characters. Recently, a benchmark dataset simulating the aforementioned input errors has been published and utilized to evaluate address parsing techniques  [111]. The third challenge is the difference in languages and address systems across regions, countries, and cultures [49]. For example, compared to United States postal addresses, the most noticeable difference in Chinese addresses is the absence of white-space delimiters. In terms of address components, the hierarchy of Chinese addresses starts from the highest administrative level, such as the State, and ends with the lowest geographic level, such as the House Number. In contrast, U.S. addresses begin with the House Number and end with the State [62].
@@
In a nutshell, the heterogeneity of geocoding input indicates that the next-generation geocoding system must be capable of handling input variations, and common input challenges such as typographic errors and semantic ambiguity need to be appropriately addressed to improve output quality of geocoding systems.
3 Geocoding Output
The key attributes of geocode outputs stem from three primary aspects: (1) geometry representations, (2) spatial accuracy, and (3) metadata description. These attributes collectively impact the usability and accuracy of geocoded data, influencing subsequent spatial analyses, interpretations, and decision-making processes.
@@ 3.1 Geometry representations
In terms of geometry, outputs from geocoding systems can take the form of points, polygons, or line segments, which represent buildings, parcels or areas, and street networks in reality  [113]. The final output geometries are determined by the underlying matching algorithms and reference datasets employed. For example, when city names are input into geocoding systems, the output can either be the polygon of the city boundary or the centroid of the city boundary. In cases where specific addresses (i.e., points) cannot be retrieved, commercial geocoding systems may revert to higher geographic levels (e.g., postal codes), resulting in boundaries of postal codes  [32]. This sensitivity to system-specific logic raises important questions and debates  [34]: Which geometry should be used as the output? Where is the optimal point to represent an entire city or postal code area? The answers to these questions often depend on the spatial resolution demands of subsequent studies. For instance, while using the centroid point of a county has been shown to be sufficient for indicating disparities across the entire country  [124], it is not accurate for analyzing the accessibility of health service providers  [70].
@@ 3.2 Spatial accuracy
Spatial accuracy in geocoding refers to the distance between geocoded locations and their corresponding ground-truth positions. It is the most critical metric for evaluating geocoding performance [36, 108, 32, 113, 120], significantly influencing subsequent spatial analyses [52, 116]. Previous studies indicate that spatial accuracy requirements and tolerances vary depending on the specific applications of geocoded data. For example, certain cancer risk assessments require spatial errors of less than 50 meters  [102], whereas, exposure classification of Toxic Release Inventory (TRI) facilities can tolerate geocoding errors up to 1500 meters  [117]. For studies linking census information with individual data, only a small portion (i.e., 5%) of geocoded data is affected by spatial error  [55], while street geocoding techniques can result in spatial errors exceeding 200 meters  [115]. Nevertheless, regardless of the analytical scale or approach, spatial accuracy significantly impacts study conclusions, especially when geocoded locations fall near aggregation or classification boundaries [89].
@@ 3.3 Metadata description
Descriptive metadata for geocoding systems are metrics that describe individual geocoded data or the performance of whole systems. The capability of metadata reporting is considered as a criteria to differentiate geocoding systems  [36]. Common individual geocode output metadata descriptions from different geocoding systems are listed in Table 2. These descriptions can be grouped as spatial and non-spatial.
@@
Spatial descriptions are generally preferred over non-spatial ones, as non-spatial descriptions lack the precision necessary for accurately ranking candidates in spatial analyses. For instance, a geocoding system may report high match rates (a non-spatial measure) yet demonstrate low spatial accuracy (a spatial measure), depending on the underlying geocoding algorithm. Consequently, relying solely on non-spatial descriptions complicates the effective comparison of geocoding systems and hinders determining their suitability for spatial analysis tasks. However, existing efforts to enhance metadata descriptions of geocodes have mostly been insufficient for assessing the quality of individual geocode outputs. These efforts have primarily focused on modeling spatial error distributions across different environments (urban vs. rural)[8], various geocoding techniques (street vs. address point)[84], diverse reference datasets (E911 vs. TIGER)[99, 114], and their impacts on different studies[78, 122]. Recent research has proposed using spatial uncertainties—the area within which the true location is likely located—as new spatial metadata to describe geocode quality [35]. However, few systems currently include this metadata due to the additional computational complexity required to calculate these uncertainties between reference data and candidate outputs [32, 35].
In summary, the quality of geocoding outputs is closely linked to the specific needs of subsequent applications, particularly concerning geometric representation and spatial accuracy. At the same time, metadata descriptions should follow standardized criteria regardless of the intended application. A standardized spatial metric (e.g., spatial uncertainty) for each geocode is highly demanded and is worthy of further exploring. Together, these three attributes discussed in this section enable diverse applications to confidently interpret and effectively utilize geocoded data.
4 Geocoding Workflow and Components
The geocoding workflow integrates multiple sophisticated components, including natural language processing, probabilistic modeling, spatial feature matching, and candidate selection methodologies [113]. In this section, we first systematically decompose this intricate workflow into distinct, manageable sub-components, followed by an in-depth exploration of each constituent element essential to the functionality of a comprehensive geocoding system.
4.1 Geocoding Workflow

### v2

@@ Abstract
Geocoding systems are widely used in both scientific research for spatial analysis and everyday life through location-based services. The quality of geocoded data significantly impacts subsequent processes and applications, underscoring the need for next-generation systems. In response to this demand, this review first characterizes the technical requirements for next-generation geocoding inputs and outputs. We then decompose the geocoding workflow into modular functional units and survey existing implementations. For each component, we identify methodological limitations, articulate domain-specific research questions and hypotheses, and outline evaluation strategies needed. Finally, we identify opportunities to improve next-generation geocoding systems in light of recent technological advances. We envision that this review provides a technical foundation and research agenda for advancing the design, assessment, and deployment of next-generation geocoding systems.
@@ 1 Introduction
The geocoding process converts human-readable location descriptions into machine-readable geographic coordinates [goldberg2007text, boscoe2008science]. The earliest known example dates back over two centuries, when a British doctor employed a dot map to visualize cholera cases [mcleod2000SnowFirstGeocoding]. In recent decades, geocoding has expanded beyond academic spatial analysis into widespread everyday applications [yin2019nlpQAframework, murray2011hybrid, rose2004historical, rushton2006geocodingforHealth, li2020modeling, yin2020probabilistic, cheng2024assessing]. This expansion has been driven by improved match rates and spatial accuracy resulting from better reference datasets, enhanced address models, and diverse data sources [amelunxen2009approach, vieira2008accuracy, zandbergen2011influence], along with rapid technological advances in web services and GPS-enabled mobile devices. Consequently, geocoding services have become readily accessible for daily use, and spatial analyses increasingly depend upon accurate geocoded data [rushton2006geocodingforHealth, murray2011hybrid].
Significant research has examined geocoding systems from both methodological and applied perspectives. Methodological studies have focused on key components of the geocoding workflow, including address parsing [Christen2005hmmAddressParserInputErrors, mokhtari2019BILSTMaddressParser, li2018hybridChineseAddressSegmentions, yin2023chatgpt], feature matching [Ranzijn2013AddressMatchingTechniques, lin2020W2VforGeocoding, Goldberg2011ImprovingMatchRateVaryingBlock, murray2011hybridGeocodesNearbyHouseNumber], interpolation [goldberg2008geocodingbest, Cayo2003, Bakshi:2004:EOS:1032222.1032251], and candidate ranking [goldberg2010candidateselectioncriteria, knoblock2017ElasticSpatialExampleGeocoding]. In parallel, comprehensive evaluation frameworks have been developed to support geocoding system design and performance assessment [goldberg2013evaluation, christen2002febrl]. Applied studies have primarily investigated the effects of geocoding quality on downstream spatial analyses [Zandbergen2012Error2SptialAnalysisSummaryofDifferentStuidies, goldberg2013evaluation]. Additionally, previous survey articles have covered geocoding from multiple angles, including sources of error and persistent challenges [goldberg2007text], implications for health-related research [rushton2006geocodingforHealth], and applications in Location-Based Services (LBS), particularly in the context of privacy protection [shin2012privacy].
@@
To inspire future advancements towards next-generation geocoding systems, this article begins by analyzing current demands for geocoding input and output as the objects to be achieved in Section 2 and Section 3, respectively. We then break down a geocoding system into key components based on their roles in the workflow, examining their existing approaches, and proposing a set of research questions and hypothesis for investigation in Section 4. Finally, we identify research directions for the development of next-generation geocoding systems in Section 5. Section 6 concludes with a summary of this survey and future envisions. This object-driven and component-based perspective differentiates our review from the existing literature.
2 Geocoding Input Demands
The input for any geocoding system is a reflection of human perception and the subjective expression of a location. For example, the description for a POI might contains place abbreviations, in contrast to the public health domain, where inputs to geocoding are typically residential addresses. These input variations place significant demands on geocoding systems to handle them effectively. To create a clear picture of the inputs that geocoding systems must handle, we examine the sources of input heterogeneity and the challenges they introduce.
Table 1 presents distinct location descriptions from the U.S. address system that have served as geocoding inputs across various studies and applications.
@@ 2 Geocoding Input Demands
Table 1: Location descriptions for different application scenarios within the U.S. address system
@@
Traffic accidents [levine1998carAccidentLocation] Distance/Topology related to intersections or landmarks 50 meters to the East of the intersection of Texas Ave. and Wellborn Rd.
Disease Outbreaks [zinszer2010diseaseOutbreak] Residential Postal Address 209 Augsburg Ct., College Station TX 77843
Crime Analysis [murray2011hybrid]
Exposure Risk Assessment [rushton2006geocodingforHealth]
Location Based Services [tiwari2011survey] Point of Interests (POI) Descriptions, Distance/Topology related to POIs Eiffel Tower, Downtown Paris
@@
Named Entity Recognition (NER) from Articles [roberts2010toponym] Administration Regions (e.g., city, state) The Statue of Liberty is in New York Harbor in New York City, in the United States.
Indoor Navigation [di2005indoor] Room Identifications + Building Identifications Room SW160, Building E
@@
These input variations can be explained by three factors. First, geocoding input describes where events or incidents often occur. Case reporting systems typically regulate the format of location descriptions during a data collection process. For instance, patients are required to provide their postal addresses upon hospital admission. Similarly, in Texas, the car crash report instructions  [^1] https://ftp.dot.state.tx.us/pub/txdot-info/trf/crash_notifications/2023/cr100-v26.1.pdf mandate that each crash location be described in relation to the nearest intersections, referenced landmarks, or highway segments. Second, geocoding input reflects human cognition of locations [hu2019semanticPlaceName]. For example, the region termed “downtown” does not physically exist on the earth but is defined by human activities, functions, and semantics [gao2017VagueRegions]. Such semantic descriptions frequently appear in daily communication, often containing abbreviations and aliases for locations [goldberg2013geocodingTechniques]. With the increasing prevalence of LBS such as ride-sharing and navigation in daily life [al2019towardsGeocodingSpatialExpressions, tiwari2011survey], geocoding inputs now range from formatted postal addresses to vernacular points of interest (POI) descriptions [melo2017SurveyOfGeocodingTextualDocuments, goldberg2013geocodingTechniques]. Third, geocoding inputs stem from how humans artificially encode spatial areas. Different encoding schemes use combinations of alphabets and digits to represent spatial partitions, which ultimately serve as geocoding inputs. The most common encoding method is postal codes, assigned based on regional delivery needs [collins1998errorsZipcode, hurley2003post]. Similarly, room identification in buildings—encoding floor levels and segmentations based on arbitrary rules—functions as an input for indoor navigation scenarios.
Table 2: Examples of address formats in the Northern and Southern hemispheres countries
Region / Country Address components
Northern Hemisphere
United States [Street] 1600 Pennsylvania Ave NW, [City] Washington, [State] DC [Postal Code] 20500, [Country] USA
France [Street] 10 Avenue des Champs-Élysées, [Postal Code] 75008 [City] Paris, [Country] France
Southern Hemisphere
Brazil [Street] Av. Paulista, 1578 – [Neighborhood] Bela Vista, [City] São Paulo – [State] SP, [Postal Code] 01310-200, [Country] Brazil
New Zealand [Street] 45 Cuba Street, [Suburb] Te Aro, [City] Wellington, [Region] Wellington Region [Postal Code] 6011, [Country] New Zealand
In addition to differences in how people describe locations, geocoding inputs also vary by language and by the address system schema used in different countries. In Table 2, we list four addresses from countries in the Northern and Southern hemisphere, respectively. The italicized text enclosed in brackets denotes the address components. For instance, the United States follows a sequence of StreetCityStatePostal CodeCountry, whereas France places the Postal Code before the City. In the Southern Hemisphere, Brazil inserts an intermediate Neighborhood component between the Street and City, while New Zealand often includes a Suburb and Region before the Postal Code. Thus, although the fundamental components of addresses are broadly similar, their order and the presence of these components largely depend on national conventions.
@@
These variations in input format and content pose significant challenges for geocoding systems. Moreover, three key linguistic complexities further complicate the geocoding process. The first challenge is the semantics and ambiguity of words. Similar to semantic issues, ambiguity is a well-known problem in geocoding and Geographic Information Retrieval (GIR) domains [hu2018GIRChallenges, goldberg2013geocodingTechniques]. The ambiguity between street direction prefixes and street names frequently appears in geocoding practice. For example, ”N Main St.” could refer to a street named ”North Main St.” or ”Main St.” with a prefix of ”North.” The second challenge is that geocoding input is error-prone. A study has categorized input errors into two types: semantic errors and syntax errors [hutchinson2010abmGeocodingInpurError]. Semantic errors are the result of the substitution of words with similar meanings. For example, a user might input ”Little River Rd” instead of ”Small Creek Rd”. Despite their similar meanings, geocoding systems struggle to distinguish them based solely on pronunciation or string similarities. Syntax errors often stem from habitual typing habits. For instance, the city name or the state name is often omitted in an address description [Christen2005hmmAddressParserInputErrors]. This omission can affect the reference feature matching process, as the same street name can be found in many cities. Besides missing address elements, typographic errors frequently occur in geocoding input and are found to be sensitive to keyboard layout [Kukich1992TypoCategroiesTypographicCognitive, Gundel1993ErrorCauseByKeyboard]. Studies have identified the most frequent typographic error patterns in geocoding input as follows [christen2002febrl, harries1999geocodingInputErrorsExample, Ranzijn2013AddressMatchingTechniques, Christen2005hmmAddressParserInputErrors, damerau1964singleErrorFreq]: (1) replacing one character, (2) deleting one character, (3) inserting one character, and (4) swapping two characters. The third challenge arises from differences in languages and address system schemas across regions, countries, and cultures [huang1996segmentationChineseLanguageCharteristizs]. The variation in the ordering of address components requires geocoding systems to go beyond a universal rule and adapt to country-specific conventions. Likewise, linguistic differences demand that geocoding systems correctly parse inputs written in different languages. For example, compared to U.S. postal addresses, Chinese addresses present the additional challenge of lacking white-space delimiters. Recently, a benchmark dataset that simulates the aforementioned input typographical errors has been published and utilized to evaluate address parsing techniques  [yin2023chatgpt]. Nonetheless, further work is still needed to create additional benchmarks that encompass a wider spectrum of input variations.
@@
In a nutshell, the heterogeneity of geocoding inputs underscores the need for next-generation geocoding systems to accommodate variations stemming from description practices, linguistic diversity, and differences in address system design. Common input challenges such as typographic errors and semantic ambiguity need to be appropriately addressed to improve output quality of geocoding systems.
3 Geocoding Output Requirements
In addition to returning spatial coordinates, geocoding systems typically produce supplementary attributes that describe the output  [goldberg2013evaluation]. Because geocoded data support critical spatial analysis workflows and applications, modern systems are expected to tailor these attributes to specific use cases [jacquez2012research]. In this section, we summarize the roles of key output attributes, assess their limitations, and propose desirable enhancements. All these analysis collectively define the requirements for next-generation geocoding systems.
The key attributes of geocoding outputs fall into three categories: (1) geometric representation, (2) spatial accuracy, and (3) metadata description. These attributes shape the usability and reliability of geocoded data.
@@ 3.1 Geometry representations
In terms of geometry, outputs from geocoding systems can take the form of points, polygons, or line segments, which represent buildings, parcels or areas, and street networks in reality  [zandbergen2008comparison]. These geometry representations are determined by input location description, the underlying matching algorithms, and reference datasets employed. For example, when city names are input into geocoding systems, the output could be either the polygon of the city boundary or the centroid of the city boundary. In cases where specific addresses (i.e., points) cannot be found in the reference datasets, commercial geocoding systems could revert to higher geographic levels (e.g., postal codes), generating the boundaries of postal codes as output  [goldberg2010candidateselectioncriteria]. This sensitivity to system-specific logic raises important questions and debates  [goldberg2007text]: Which geometry should be used as the output? Where is the optimal point to represent an entire city or postal code area? The answers to these questions often depend on the spatial resolution demands of subsequent studies. For instance, while using the centroid point of a county has been shown to be sufficient to represent disparities across a country  [zou2018TwitterDisaster], it is not accurate enough when analyzing the accessibility of health service providers  [McDonald2017HealthAccessibilityGeocoding]. Therefore, it is crucial that any decision regarding geometric representation is clearly reflected in the geocoding output so that end users can make informed decisions about how to use the resulting geocodes. Yet, few existing systems provide this information.
@@ 3.2 Spatial accuracy
Spatial accuracy in geocoding refers to the distance between geocoded locations and their corresponding ground-truth positions. It is the most critical metric for evaluating geocoding performance [goldberg2013evaluation, yin2019DLRooftopGeocoding, goldberg2010candidateselectioncriteria, zandbergen2008comparison, zhang2024mining], significantly influencing subsequent spatial analyses [jacquez2012research, Zandbergen2012Error2SptialAnalysisSummaryofDifferentStuidies]. Previous studies indicate that spatial accuracy requirements and tolerances vary depending on the specific applications of geocoded data. For example, certain cancer risk assessments require spatial errors of less than 50 meters  [washburn199geocodedError50meters], whereas exposure classification of Toxic Release Inventory (TRI) facilities can tolerate geocoding errors up to 1500 meters  [Zhan2006geocodingTWI1500meters]. For studies linking census information with individual data, only a small portion (i.e., 5%) of geocoded data is affected by spatial error  [Kravets2007PointInPolygonError], while street geocoding techniques can result in spatial errors exceeding 200 meters  [zandbergen2007exposure250MBiasResults]. Nevertheless, regardless of the analytical scale or approach, spatial accuracy significantly impacts study conclusions, particularly when geocoded locations fall near aggregation or classification boundaries [schootman2005geocodedPointInPolygon]. For example, when a geocode falls near the boundary of two house lots and its position is uncertain, providing information about the likelihood of association with each lot can be highly valuable. Thus, we argue that geocoding systems should provide a means to quantify and indicate the spatial relationship between output geocodes and nearby features. We will elaborate on this requirement in Section 3.3.
@@ 3.3 Metadata description
Descriptive metadata for geocoding systems are metrics that describe the quality of each individual or the entire geocoded output collection. The capability of reporting metadata is considered as a crucial criteria to differentiate geocoding systems  [goldberg2013evaluation]. Common geocode output metadata descriptions from different geocoding systems are listed in Table 3. These descriptions can be grouped as spatial and non-spatial.
@@
Generally, spatial descriptions are preferred over non-spatial ones, non-spatial descriptions lack the information needed to accurately evaluate the quality of geocodes. For example, given a set of geocoding inputs, depending on the underlying algorithm, a system may report high match rates (a non-spatial metric) by reverting a less spatial accuracy level to find a match. Consequently, relying solely on the match rate without considering the spatial accuracy can negatively impact the potential usage of output geocodes, particularly when one needs to select an appropriate system for data processing.
To date, research has primarily focused on evaluating collections of geocoded outputs rather than individual results. This includes modeling spatial error across different environments [Cayo2003ResidualUrbanRural], geocoding techniques [roongpiboonsopit2010quality], and reference datasets [vieira2008accuracy, zandbergen2011influence]. Yet these models offer limited insight into the quality of individual geocodes. Recent research has proposed using spatial uncertainties—the area within which the true location is likely located—as new spatial metadata to describe geocode quality [goldberg2010towardQuantitativeGeocodeAccuracy]. As aforementioned, when output geocodes fall close to the boundaries of areas such as parcels or counties, quantifying the likelihood that the geocode belongs to each area is of importance. However, few systems currently include this metadata, as it requires integrating a quantification algorithm into existing geocoding workflow to calculate these uncertainties between reference data and candidate outputs [goldberg2010candidateselectioncriteria, goldberg2010towardQuantitativeGeocodeAccuracy].
Given the facts presented above, we can conclude that the choice of geometric representation and spatial accuracy level should be closely tied to the specific needs of subsequent application scenarios. In addition, the spatial metadata associated with individual geocodes—or with a set of geocodes—can be critical for downstream tasks, helping users determine whether the outputs are suitable and what potential biases may arise from their usage. Accordingly, we propose the following new requirements that next-generation geocoding systems should meet:
  • providing a means to illustrate the rationale behind choosing one geometric representation for output over others, especially when spatial accuracy is sacrificed when attempting to identify a geocoded.
  • reporting the necessary spatial uncertainty for each individual geocode in relation to nearby features, particularly when the geocode lies close to boundaries or other spatially adjacent features.
4 The State of Geocoding Components
Having defined the standards for geocoding inputs and outputs, we now turn to the geocoding process, which accepts these inputs and generates the corresponding outputs. Geocoding is a complex workflow involving natural language processing, probabilistic retrieval, spatial feature matching, and candidate selection [zandbergen2008comparison]. To identify solutions that meet the high standards defined above, it is essential to examine each component of this workflow in greater depth and to pinpoint the gaps that may still remain. Thus, in this section, we first decompose the geocoding process into distinct and manageable components. We then describe the functionality and current state of each component in detail. Finally, we outline the challenges and the future research directions for each component when building the next-generation geocoding systems.
4.1 Geocoding Workflow Decomposition